OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, 6) and has a slope of –1?

OpenStudy (anonymous):

– y = –4 x + y = 4 x – y = 4 x + y = –4

OpenStudy (anonymous):

y=-x+q with undefined q, put then the x and y of the point here.... 6=2+q so q=4 finally the eqution will be y=-x+4

OpenStudy (anonymous):

xD in this case solution is the second one ;)

OpenStudy (anonymous):

not coinvinced yet?